Solve the equation.

y 2 - 15 y + 44 = 0

Answer by checkley71(8403)

Y^2-15Y+44=0
(Y-11)(Y-4)=0
Y-11=0
Y=11 ANSWER.
Y-4=0
Y=4 ANSWER.